OpenStudy (anonymous):

A bag of fruit contains 3 apples and 2 oranges and 1 banana and 4 pears.Gerald will randomly selected two pieces of fruit one at a time from the bag and not put is back. What is the probability that the first piece of fruit Gerald selects will be a banana and the second piece of fruit will be a pear?? F. 1/25 G. 2/45 H. 1/2 J. 4/9 Please help I'll give a medal!!!

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

First find tjhe total amount of fruit.

OpenStudy (anonymous):

10

OpenStudy (studygurl14):

Correct. Now find the amount of bananas.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

In Circle p, what is the measure of ADB?

OpenStudy (studygurl14):

Correct. now find the amount of pears @islander691 Please do not spam. Do not post your question in someone else's thread

OpenStudy (anonymous):

lol. srry i jus need help.

OpenStudy (anonymous):

4

OpenStudy (studygurl14):

Tag me to your question. @islander691

OpenStudy (anonymous):

aight

OpenStudy (anonymous):

In Circle p, what is the measure of ADB?

OpenStudy (studygurl14):

correct. now you need to multiply the probailities of choosing a banana and choosing a pear together. So do this: \(\Large \frac{1}{10}\times\frac{4}{10}\)

OpenStudy (studygurl14):

I said tag me to your question @islander691 Open your question in your own post, and then tag me by typing @StudyGurl14

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

4/10 or 4/20

OpenStudy (anonymous):

....

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I already got the answer....long time ago..but still thank u